Solve the systems 16= x^2 + y^2, (y-2)^2 + (x-1)^2 =9 involving quadratic equations using any appropriate method.
Answer by Alan3354(69443)

Solve the systems 16= x^2 + y^2, (y-2)^2 + (x-1)^2 =9 involving quadratic equations using any appropriate method.
-------------------------
16= x^2 + y^2
(y-2)^2 + (x-1)^2 =9
---
16= x^2 + y^2
(y-2)^2 + (x-1)^2 =9
---
x^2 + y^2 = 16
x^2 - 2x + y^2 - 4y = 4
---
x^2 - 2x + y^2 - 4y = 4
x^2 + y^2 = 16
------------------------------------------- Subtract
-2x - 4y = -12
x + 2y = 6
x = 6 - 2y
=====================
x^2 + y^2 = 16
(6-2y)^2 + y^2 = 16
5y^2 - 24y - 16 = 0
---
Solve for y --> 2 solutions
sub each and find 2 for x.